# FedDUAL: A Dual-Strategy with Adaptive Loss and Dynamic Aggregation for Mitigating Data Heterogeneity in Federated Learning

### v1

@@ 4.1 Experimental Setup
To assess the effectiveness of the proposed FedDUAL approach, we conducted extensive experiments using three widely recognized classification benchmarks: CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky, 2009), and FMNIST (Xiao et al., 2017). To simulate real-world non-IID data distributions, we employed a client-wise partitioning strategy based on the Dirichlet distribution (Hsu et al., 2019). This distribution is governed by a concentration parameter $\alpha$, which controls the degree of data heterogeneity among clients. Lower $\alpha$ values result in more skewed data distributions, closely mimicking uneven data partitions. In all experiments, we set $\alpha$ = 0.01 to simulate severe data heterogeneity, closely approximating real-world conditions. Throughout the communication rounds, each client retains a fixed local data partition. To evaluate the global model’s classification performance, we use a separate test dataset maintained at the server, which remains unseen during training. For our experiments, we used LeNet (LeCun et al., 1998) for FMNIST dataset and a pre-trained VGG16 (Simonyan & Zisserman, 2015) for CIFAR-10 and CIFAR-100 dataset, following the methodology outlined in (Hu et al., 2024). We applied the proposed dynamic aggregation mechanism only to the last two layers of these models. Our setup involved $100$ clients, with 10% randomly sampled per communication round, and a batch size of $32$. Each client performed three local epochs of model updates. We have computed each result three times with different seed values and reported the mean value with standard deviation. To determine the optimal client learning rate for each experiment, we conducted a grid search over ${0.05,0.01,0.2,0.3}$. For the baseline FedProx, we tested proximal values of 0.001, 0.1, 0.4, 0.7 to find the optimal setting, and for FedNova, we evaluated proximal SGD values from 0.001, 0.003, 0.05, 0.1, following the recommendations in Li et al. (2024). Across all experiments, we used the Adam optimizer for consistency. We have run each algorithm three times and reported the average outcome. The experimental setup utilized an NVIDIA Quadro RTX 4000 GPU boasting 40GB of memory. The implementation was crafted using Python  [^1] https://www.python.org/, leveraging the TensorFlow framework  [^2] https://www.tensorflow.org/ utilizing Windows 11.
@@ 4.2 Comparison with the State-of-the-art Methods
We also provide a comparative review of learning process, as shown in Fig. 3, where we present the learning curves of our method with the baselines. The results demonstrate a consistent pattern: our approach exhibits accelerated learning and achieves superior accuracy across all tested scenarios. While the communication rounds vary by dataset, we observe that model performance reaches a saturation point by the conclusion of these rounds. Notably, our method not only attains a more robust final model but also displays markedly faster convergence across all datasets examined. This effectiveness is further highlighted in Fig. 4, where it consistently reaches target accuracy with far fewer communication rounds compared to baseline approaches.
@@ 4.3 Validation of the Motivation
Table 2: Ablation study results for the proposed FedDUAL method.
Adaptive Loss Dynamic Aggregation Accuracy (%)
80.70 $\pm$ 0.22
@@
5 Ablation Study
In our ablation study, we performed all experiments on the FMNIST dataset with $\alpha=0.01$. The study comprised four types of experiments: (1) performance analysis of the individual modules, (2) assessment of the impact of dynamic aggregation across different neural network layers, (3) hyperparameter analysis, and (4) evaluation of various levels of data heterogeneity. The results of the first and second experiments are detailed in the following subsections, while the third and fourth experiments are presented in Sections F of the Appendix.
5.0.1 Performance Analysis of Individual Modules.
To demonstrate the significance of our proposed adaptive loss function and dynamic aggregation technique, we conducted three distinct experiments, with results summarized in Table 2. The first experiment utilized only the adaptive loss function, paired with simple server-side aggregation. The second implemented our dynamic aggregation technique at the server, while retaining the conventional cross-entropy loss function locally. Finally, the third experiment combined both proposed methods: the adaptive loss function and dynamic aggregation technique. As evidenced by Table 2, the integration of both proposed approaches in the third experiment yielded the highest accuracy, highlighting the impact of our dual strategy on model performance. The learning curves for these experiments are illustrated in Fig. 7 of the Appendix.
@@ 5 Ablation Study
5.0.2 Impact of Dynamic Aggregation on Different Network Layers.
6 Limitation
The proposed FedDUAL framework tackles heterogeneous data challenges in federated training, outperforming baselines through an adaptive client-side loss function and dynamic server-side aggregation via the Wasserstein Barycenter. While the proposed method significantly enhance performance and accelerate convergence, they introduce additional computational overhead at the server due to the iterative calculation of the Wasserstein Barycenter. Conversely, the adaptive loss function on the client side is computationally efficient, incurring minimal cost. Future research will focus on reducing the computational burden of the Wasserstein Barycenter calculation by developing more efficient algorithms, aiming to maintain or even improve the performance of the proposed framework.
@@ 7 Conclusion
Appendix A Appendix
Appendix B Convergence Proof
B.1 Preliminaries
B.1.1 Notation
Let $\mathcal{W}$ denote the parameter space, $f_{i}:\mathcal{W}\rightarrow\mathbb{R}$ be the loss function for the $i^{th}$ client, and $F:\mathcal{W}\rightarrow\mathbb{R}$ be the global objective function.
B.2 Assumptions
1. Lipschitz Smoothness: For all $i$, $f_{i}$ is L-smooth such that it follows Eq. 14.
@@ B.2 Assumptions
$\|\nabla f_{i}(w)-\nabla f_{i}(w^{\prime})\|\leq L\|w-w^{\prime}\|,\forall w,w^{\prime}\in\mathcal{W}$ (14)
2. Bounded Variance: The variance of stochastic gradients is bounded by Eq. 15.

### v2

@@ 5.1 Experimental Setup
To assess the effectiveness of the proposed FedDUAL approach, we conducted extensive experiments using three widely recognized classification benchmarks: CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky, 2009), and FMNIST (Xiao et al., 2017). To simulate real-world non-IID data distributions, we employed a client-wise partitioning strategy based on the Dirichlet distribution (Hsu et al., 2019). This distribution is governed by a concentration parameter $\alpha$, which controls the degree of data heterogeneity among clients. Lower $\alpha$ values result in more skewed data distributions, closely mimicking uneven data partitions. In all experiments, we set $\alpha$ = 0.01 to simulate severe data heterogeneity, closely approximating real-world conditions. Throughout the communication rounds, each client retains a fixed local data partition. To evaluate the global model’s classification performance, we use a separate test dataset maintained at the server, which remains unseen during training. For our experiments, we used LeNet (LeCun et al., 1998) for FMNIST dataset and a pre-trained VGG16 (Simonyan and Zisserman, 2015) for CIFAR-10 and CIFAR-100 dataset, following the methodology outlined in (Hu et al., 2024). We applied the proposed dynamic aggregation mechanism only to the last two layers of these models. Our setup involved $100$ clients, with 10% randomly sampled per communication round, and a batch size of $32$. Each client performed three local epochs of model updates. We have computed each result three times with different seed values and reported the mean value with standard deviation. To determine the optimal client learning rate for each experiment, we conducted a grid search over 0.05,0.01,0.2,0.3{0.05,0.01,0.2,0.3}. For the baseline FedProx, we tested proximal values of 0.001, 0.1, 0.4, 0.7 to find the optimal setting, and for FedNova, we evaluated proximal SGD values from 0.001, 0.003, 0.05, 0.1, following the recommendations in Li et al. (2024). Across all experiments, we used the Adam optimizer for consistency. We have run each algorithm three times and reported the average outcome. The experimental setup utilized an NVIDIA Quadro RTX 4000 GPU boasting 40GB of memory. The implementation was crafted using Python  [^1] https://www.python.org/, leveraging the TensorFlow framework  [^2] https://www.tensorflow.org/ utilizing Windows 11.
@@ 5.2 Comparison with the State-of-the-art Methods
Figure 3 compares the learning curves of our method with baselines, while Fig. 10 in the Appendix includes the corresponding curves with error bars. Across all datasets, our method consistently converges faster and achieves higher final accuracy. Although the number of communication rounds varies by dataset, performance generally saturates by the final round. Notably, our method not only attains a more robust final model but also displays markedly faster convergence across all datasets examined. This effectiveness is further highlighted in Fig. 4, where it consistently reaches target accuracy with far fewer communication rounds compared to baseline approaches. However, the proposed method incurs a higher computation cost due to Wasserstein barycenter based aggregation at the server side. We have provided a thorough computational complexity analysis and runtime analysis in the Section C and the Section D of the appendix respectively.
@@ 5.3 Validation of the Motivation
6 Ablation Study
In our ablation study, we performed all experiments on the FMNIST dataset with $\alpha=0.01$. The study comprised four types of experiments: (1) performance analysis of the individual modules, (2) assessment of the impact of dynamic aggregation across different neural network layers, (3) hyperparameter analysis, and (4) evaluation of various levels of data heterogeneity.
6.0.1 Performance Analysis of Individual Modules
To assess the effectiveness of the proposed adaptive loss and dynamic aggregation techniques, we conducted three ablation experiments across FMNIST, CIFAR-10, and CIFAR-100. The results for FMNIST are shown in Table 2. In the first experiment, we employed only the adaptive loss alongside standard server-side aggregation. Notably, this configuration underperforms FedAvg across all datasets, indicating that adaptive loss alone cannot effectively address data heterogeneity—likely due to its limited capacity to improve generalization despite fostering local-global alignment. The second experiment implemented our dynamic aggregation technique at the server, while retaining the conventional cross-entropy loss function locally. Finally, the third experiment combined both proposed methods: the adaptive loss function and the dynamic aggregation technique. As evidenced by Table 2, the integration of both proposed approaches in the third experiment yielded the highest accuracy, highlighting the impact of our dual strategy on model performance. The learning curves for these experiments using FMNIST dataset are illustrated in Fig. 7 of the Appendix.
Table 2: Ablation study of FedDUAL across different datasets.
Adaptive Loss Dynamic Agg. Dataset Acc. (%)
FMNIST 80.70 $\pm$ 0.22
@@ 6.0.1 Performance Analysis of Individual Modules
CIFAR10 41.05 $\pm$ 0.12
46.50 $\pm$ 0.15
48.70 $\pm$ 0.20
CIFAR100 25.05 $\pm$ 0.11
27.01 $\pm$ 0.17
29.15 $\pm$ 0.24
6.0.2 Impact of Dynamic Aggregation on Different Network Layers
6.0.3 Hyperparameter Analysis
In the proposed architecture, there are two key hyperparameters to consider: the scaling factor ($\gamma$) and the number of iterations used to compute the Wasserstein Barycenter. The proposed FedDUAL approach utilizes dynamic server-side aggregation by applying the Wasserstein Barycenter concept to combine the weights of the final layers from local models. This iterative process involves a small positive constant ($\epsilon$) to determine the scaling factor ($\gamma$). To optimize performance, we conducted two sets of experiments on the FMNIST dataset with $\alpha=0.01$, each exploring a range of values for these crucial hyperparameters. The hyperparameter $\epsilon$ influences the sensitivity of the barycenter calculation to variations in Wasserstein distance. A smaller $\epsilon$ makes the barycenter more responsive to differences in Wasserstein distance, while a larger $\epsilon$ diminishes this sensitivity. This impacts how the barycenter integrates each distribution according to its distance from the current estimate. During the iterative update of the barycenter, $\epsilon$ affects the scaling factor $\gamma$ applied to each distribution. An excessively small $\epsilon$ can result in slow or potentially non-existent convergence due to minimal scaling factor, whereas a too-large $\epsilon$ may cause oversmoothing, reducing the barycenter’s effectiveness in accurately representing the distributions. For this setting we have fixed the number of iterations to compute Wasserstein Barycenter as 150. Figure 8 in the Appendix shows test accuracy across different $\epsilon$ values, indicating that larger $\epsilon$ can degrade performance or hinder convergence. Figure 9 in the Appendix presents the corresponding learning curves for these settings. The number of iterations in the Wasserstein Barycenter function is another critical hyperparameter that affects both the accuracy and efficiency of the barycenter computation. Generally, more iterations enhance convergence and accuracy, ensuring that the barycenter more closely approximates the optimal value. However, increasing the number of iterations also prolongs computation time, necessitating a balance between accuracy and efficiency. Finding the optimal number of iterations involves a trade-off: too few iterations may result in suboptimal outcomes, while too many can yield diminishing returns in accuracy. To achieve the best performance, begin with a reasonable default value, monitor convergence by observing changes in the barycenter, and adjust iteratively based on empirical results and available computational resources. For this setting, we fixed the epsilonepsilon value as 0.0001, which yields the highest results in previous experiment. Figure 11 illustrates the test accuracy for different values of iterations to calculate Wasserstein Barycenter, suggesting that larger iterations may adversely affect performance. Figure 12 presents the corresponding learning curves for these settings. From both experiments, we observe that the highest performance is achieved with $\epsilon=0.00001$ and 150 iterations. Therefore, to optimize performance, it is advisable to set $\epsilon$ to a smaller value while keeping the number of iterations between 100 and 150.
@@ 6.0.3 Hyperparameter Analysis
6.0.4 Experiment on Different Level of Data Heterogeneity
Figure 13 illustrates the accuracy of the proposed method and various baselines across different levels of data heterogeneity on the FMNIST dataset. In this context, heterogeneity is quantified by $\alpha$, with lower values indicating greater data heterogeneity. The results show that as $\alpha$ decreases, the test accuracy for all models increases, because data heterogeneity among clients is decreased. Remarkably, the proposed method consistently achieves the highest test accuracy and exhibits the slowest performance decline compared to other algorithms, demonstrating superior performance of the proposed method on varying degrees of non-IID data partitioning. The learning curve is presented in Fig. 14.
@@ 7 Conclusion
Appendix
Appendix A Convergence Proof
Before presenting the main convergence theorems, we establish assumptions and several key lemmas.
Assumptions
1. L-Smoothness: Each local loss function $f_{k}(\theta)$ is $L$-smooth:
@@
$\lVert\nabla f_{k}(\theta)-\nabla f_{k}(\theta^{\prime})\rVert\leq L\lVert\theta-\theta^{\prime}\rVert,\quad\forall\theta,\theta^{\prime}\in\mathcal{W},\forall k$ (17)
2. Unbiased Stochastic Gradients: For any client $k$ and parameter $\theta$, the stochastic gradient is unbiased:
$\mathbb{E}_{\xi\sim\mathcal{D}_{k}}[\nabla\ell(\theta;\xi)]=\nabla f_{k}(\theta).$ (18)
3. Bounded Variance: The variance of stochastic gradients is bounded: